# Towards Definition of Higher Order Causality in Complex Systems

### v1

@@ Abstract
The description of the dynamics of complex systems, in particular the capture of the interaction structure and causal relationships between elements of the system, is one of the central questions of interdisciplinary research. While the characterization of pairwise causal interactions is a relatively ripe field with established theoretical concepts and the current focus is on technical issues of their efficient estimation, it turns out that the standard concepts such as Granger causality or transfer entropy may not faithfully reflect possible synergies or interactions of higher orders, phenomena highly relevant for many real-world complex systems. In this paper, we propose a generalization and refinement of the information-theoretic approach to causal inference, enabling the description of truly multivariate, rather than multiple pairwise, causal interactions, and moving thus from causal networks to causal hypernetworks. In particular, while keeping the ability to control for mediating variables or common causes, in case of purely synergetic interactions such as the exclusive disjunction, it ascribes the causal role to the multivariate causal set but not to individual inputs, distinguishing it thus from the case of e.g. two additive univariate causes. We demonstrate this concept by application to illustrative theoretical examples as well as a biophysically realistic simulation of biological neuronal dynamics recently reported to employ synergetic computations.
@@ I Introduction
A common approach is the Granger causality [2] - a concept based on two principles: the cause happens before its effect, and the cause carries some additional information about the effect (not included in the ’rest of the universe’). A nonlinear generalization of this concept - transfer entropy [3, 4] - is based on the same principles. While Granger causality is typically cast in the framework based on prediction via linear vector autoregressive processes, transfer entropy is an information-theoretic measure aiming to capture time-directed information transfer of arbitrary functional form. Indeed, for Gaussian processes, the two concepts are equivalent [5]. It is useful to conceptualize Granger causality as a fundamental refinement of the notoriously naive concept of causality as correlation by adding the requirement of the candidate cause holding correlation (prediction/information) on top of that included in the target past and the rest of the universe. This key refinement allows ascribing the causal effect more conservatively, in particular controlling for common sources or mediating variables. In this work, we suggest yet another refinement that would avoid ascribing causal status to variables, that only contribute through multivariate interactions, while keeping the original safeguards of Granger’s approach.
Another prominent approach to causality, developed by Judea Pearl [6], assumes that so-called interventions can be carried out, making it a more difficult concept to apply due to the need to be able to carry out the interventions and assess their effects (analytically or experimentally). On the other hand, unlike Granger causality or transfer entropy, Pearl’s framework does not assume the temporal organization of the investigated data, allowing thus causal inference even without sampling time series with proper temporal resolution. Nevertheless, the approaches are to a large extent related and, in particular, describe the causal structure of the investigated system as a directed graph between the variables.
@@
We suggest that it is useful to generalize this view to consider representing causal structure by hypergraphs rather than graphs/networks. Let us consider a pristine example of purely higher order interaction between source variables in causing the target variables: two candidate source variables $X_{1},X_{2}\stackrel{{\scriptstyle iid}}{{\sim}}Be(0.5)$ and a target variable $Y$ defined as $Y=XOR(X_{1},X_{2}),$ or in the context of time series $Y(t+1)=XOR(X_{1}(t),X_{2}(t)).$ Note that logical XOR is zero if the inputs are equal; otherwise, it is one. In this system, pairwise mutual information $I(X_{1},Y)$ between $X_{1}$ and $Y$ is equal to zero, and so is $I(X_{2},Y)$. In other words, neither of the candidate source variables carries (on their own) information about the target variable $Y$; they play a role, but only together. The XOR function provides a pristine example of purely higher order interaction between source variables in causing the target variables. Nevertheless, Granger would ascribe causal status to both $X_{1}$ and $X_{2}$ (by conditioning each one on the other). One may, however, argue that such representation by a graph with the causal links (from $X_{1}$ to $Y$ and from $X_{2}$ to $Y$) is obfuscating something important, as it does not distinguish this system from, say, $Y=X_{1}+X_{2}$, where each variable indeed carries information about $Y$ on its own. The intuition suggests one should not exactly speak about the causal effect between $X_{1}$ (or $X_{2}$) and $Y$: they are even independent. At the same time, representing the system by an empty graph and denying causal status to both variables seems intuitively incorrect as well, as that would collide with the case of completely unpredictable $Y$. Thus, maybe most intuitively, as only the multivariate information $I(\{X_{1},X_{2}\},Y)$ is non-zero (in particular 1 bit), we suggest to state that there is no causal link, but there is a causal hyperlink from the set $\{X_{1},X_{2}\}$ to $Y$. Overall, the system should not be represented by a directed graph but by a directed hypergraph.
@@
While the XOR function may seem artificial, higher-order interactions are commonly discussed in the context of modelling complex systems in physics [7], neuroscience [8, 9], ecology [10], social sciences [11] and many more.
Despite this challenge is increasingly recognized in the causal inference community [12, 13, 14], formal refinement of causality methods such as transfer entropy in this direction is still missing – a gap we aim to fill.
@@ II Transfer entropy & Problem statement
In the following text, we assume that for each target variable $Y$, a set of candidate source variables $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ is known a priori. In practice, such assumption is justified either by Granger causality-like reasoning: the candidate source variables are all those temporally preceding the target variable, or by (additional) theoretical arguments. The transfer entropy TE (the positivity of which defines the presence of causal effect from $X_{i}$ to $Y$) is then defined as the conditional mutual information between $X_{i}$ and $Y$ conditioned on all other potential sources (i.e. $\mathbf{X}\smallsetminus\{X_{i}\}$):
@@
As already mentioned, we aim to generalize the transfer entropy concept to more faithfully represent causal structure in the presence of higher-order causal interactions. A natural generalization is to define existence of a causal effect of a set $\mathbf{X}_{I}=\{X_{i_{1}},\ldots,X_{i_{k}}\}$ on target $Y$ as
@@
What would such generalization of TE (or Pearl’s direct causal effect inference) conclude? First, it infers a causal link from the set $\{X_{1},X_{2}\}$: $I(\{X_{1},X_{2}\},Y)=1$ bit. However, it also infers causal link from $X_{1}$ ($X_{2}$), as
@@
As a side note, in practice, these individual links might remain undetected, as commonly used causal network inference algorithms progress iteratively from testing univariate predictors and may stop before discovering such conditional dependence – see [13] for discussion, and [14] for an example of an updated (yet, therefore, slower) algorithm more robust in this regard. Also note that the XOR example is straightforwardly generalizable to the k-variable higher order interaction by considering $Y$ as a sum of variables $X_{1},X_{2},\ldots X_{k-1}\stackrel{{\scriptstyle iid}}{{\sim}}Be(0.5)$ modulo 2. Again, even the knowledge of all but one of the source variables provides no information about $Y$ without knowing the last one; it is the interaction of all $k-1$ sources that provides the target predictability.
Setting aside the implementation details, even such multivariate TE definition does not distinguish the distinctive synergistic causal effect in the XOR example, compared to, e.g. the case of summing two independent normally distributed variables. Two issues arise: first, that TE concludes that $X_{1}$ has a ’causal effect’ on $Y$ although $X_{1}$ does not hold any information about $Y,$ might be considered superfluous, as the information ’belongs’ to the pair (which multivariate TE correctly marks). On the other hand, in the case of the linear sum function, as the pair does not carry information on top of each variable, one might not want to mark the pair as causal per se, but only the individual variables.
@@ III Causal effect
The system is constructed such that $X_{1}$ does not directly affect $Y$ by itself, but it affects $Y$ indirectly through mediating variables $X_{p+1},\ldots,X_{n}.$ Also $X_{1}$ affects $Y$ through pairwise interactions: each pair $\{X_{1},X_{i}\}$ for $i\in\{2,\ldots,p\}$ affects $Y$. Note that both the mutual information $I(X_{1},Y)$ and the fully conditional mutual information (transfer entropy) $I\left(\left(X_{1},Y\right)|\mathbf{X}\smallsetminus\{X_{1}\}\right)$ are greater than zero, suggesting thus a causal effect of $X_{1}$ on $Y$. In fact, all of the conditional mutual informations $I(X_{1},Y|\mathbf{S})$ are positive, except $I\left(X_{1},Y|X_{p+1},\ldots,X_{n}\right)$.
@@
Definition 1.
Let $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ be a set of candidate source variables and $Y$ a target variable. There is a causal effect from set $\{X_{i_{1}},\ldots,X_{i_{k}}\}$ to $Y$ if and only if
@@
Notably, the definition of the causal effect of a set of variables typically enforces the causal influence of an arbitrary set containing a variable with a causal effect. Consider, for example, a simple linear system where the variable $Y$ is causally affected just by variable $X_{1}$ plus some noise: $Y=X_{1}+\mathcal{E}_{Y}$, and in the system, there are also other variables $\{X_{2},\ldots,X_{n}\}$ independent on $Y$ and $X_{1}$. Then, all sets containing $X_{1}$ are also causally affecting $Y$. Note that similar behaviour (supersets of causes defined as causes irrespective of further value in prediction) also holds for (a multivariate variant of) transfer entropy or Pearl’s causality. In other words, although the Definition 1 of causal effect avoids attributing higher-order causal effects to individual variables, it still does not by itself distinguish multivariate causes that are trivial in the sense of inherited from its subsets from nontrivial higher-order multivariate causal effects.
We suggest one may add this distinction of a unique higher-order causal effect by interpreting differently the observation of multivariate causes in the presence/absence of the causal effects of the subsets. Further, to distinguish purely additive from synergistic multivariate causality, one may quantify the higher order causal information not induced trivially by the subsets; suitable approaches [15, 8, 16] are the subject of further research.
@@
Returning to the XOR example, note that the resulting causal graph of the system also depends on the distribution of the variables $X_{1}$ and $X_{2}.$ Consider three different distributions, see Table 1. In all cases ($p_{1},p_{2},p_{3}$), $X_{1}$ and $X_{2}$ are independent variables with Bernoulli distribution.
$X_{1}$ $X_{2}$ $Y=XOR\left(X_{1},X_{2}\right)$ $p_{1}$ $p_{2}$ $p_{3}$
0 0 0 0.25 0.10 0.06
0 1 1 0.25 0.40 0.24
1 0 1 0.25 0.10 0.14
1 1 0 0.25 0.40 0.56
Table 1: Function $XOR(X_{1},X_{2}),$ where $X_{1}$ and $X_{2}$ are independent random variables with distributions: $i)$ $p_{1}:\ X_{1},X_{2}\sim Be(0.5),$ $ii)$ $p_{2}:\ X_{1}\sim Be(0.5),\ X_{2}\sim Be(0.8)$ $iii)$ $p_{3}:\ X_{1}\sim Be(0.7),\ X_{2}\sim Be(0.8).$
@@
We have already discussed the first case with uniformly distributed probabilities. In the second case, $X_{1}$ is uniformly distributed, but $X_{2}\sim Be(0.8).$ While $I(X_{2},Y)$ remains zero, $I(X_{1},Y)$ is positive, as is $I(X_{1},Y|X_{2}).$ Variable $X_{1}$ thus has a causal effect on $Y$ only due to the change in the distribution $X_{2}.$ The further the variable $X_{2}$ deviates from the uniform distribution, the more information $X_{1}$ has about $Y$ and the unique contribution of the pair $\{X_{1},X_{2}\}$ decreases. In the last case, where both variables deviate from the uniform distribution, both $X_{1}$ and $X_{2}$ have a causal effect on $Y$, and they also have a unique contribution as a pair $\{X_{1},X_{2}\}.$

### v2

# Higher order definition of causality by optimally conditioned transfer entropy

@@ Abstract
The description of the dynamics of complex systems, in particular the capture of the interaction structure and causal relationships between elements of the system, is one of the central questions of interdisciplinary research. While the characterization of pairwise causal interactions is a relatively ripe field with established theoretical concepts and the current focus is on technical issues of their efficient estimation, it turns out that the standard concepts such as Granger causality or transfer entropy may not faithfully reflect possible synergies or interactions of higher orders, phenomena highly relevant for many real-world complex systems. In this paper, we propose a generalization and refinement of the information-theoretic approach to causal inference, enabling the description of truly multivariate, rather than multiple pairwise, causal interactions, and moving thus from causal networks to causal hypernetworks. In particular, while keeping the ability to control for mediating variables or common causes, in case of purely synergistic interactions such as the exclusive disjunction, it ascribes the causal role to the multivariate causal set but not to individual inputs, distinguishing it thus from the case of e.g. two additive univariate causes. We demonstrate this concept by application to illustrative theoretical examples as well as a biophysically realistic simulation of biological neuronal dynamics recently reported to employ synergistic computations.
@@ I Introduction
A common approach is the Granger causality [2] - a concept based on two principles: the cause happens before its effect, and the cause carries some additional information about the effect (not included in the ’rest of the universe’). A nonlinear generalization of this concept - transfer entropy [3, 4] - is based on the same principles. While Granger causality is typically cast in the framework based on prediction via linear vector autoregressive processes, transfer entropy is an information-theoretic measure aiming to capture time-directed information transfer of arbitrary functional form. Indeed, for Gaussian processes, the two concepts are equivalent [5].
In this work, we are proposing an extension to the framework of Granger causality (seen as reduction in variance) or Transfer Entropy (seen as reduction in surprise). To understand the nature and motivation of the extension, it is elucidating to start by conceptualizing Granger causality as a fundamental refinement of the notoriously naive concept of causality as correlation by adding the requirement of the candidate cause holding correlation (prediction/information) on top of that included in the target past and the rest of the universe. Note that analogously Transfer Entropy is a refinement of Mutual information by conditioning on target past (and potentially other sources). This key refinement allows ascribing the causal effect more conservatively, in particular controlling for common sources or mediating variables, however is not free of interpretation problems. To alleviate these, we suggest yet another refinement that would in particular avoid ascribing causal status to variables, that only contribute through multivariate interactions. We aim to do this while keeping the original safeguards (against false causal inferences) of Granger’s approach .
@@
Let us consider a pristine example of purely higher order interaction between source variables in causing the target variables: two candidate source variables $X_{1},X_{2}\stackrel{{\scriptstyle iid}}{{\sim}}Be(0.5)$, i.e. independent, each with a uniform Bernoulli distribution with probability $p=0.5$; and a target variable $Y$ defined as $Y=XOR(X_{1},X_{2}),$ or in the context of time series $Y(t+1)=XOR(X_{1}(t),X_{2}(t)).$ Note that logical XOR is zero if the inputs are equal; otherwise, it is one. In this system, pairwise mutual information $I(X_{1},Y)$ between $X_{1}$ and $Y$ is equal to zero, and so is $I(X_{2},Y)$. In other words, neither of the candidate source variables carries (on their own) information about the target variable $Y$; they play a role, but only together. Nevertheless transfer entropy ascribes causal status to both $X_{1}$ and $X_{2}$ (by conditioning each one on the other). One may, however, argue that such representation by a graph with the causal links (from $X_{1}$ to $Y$ and from $X_{2}$ to $Y$) is obfuscating something important, as it does not distinguish this system from, say, $Y=X_{1}+X_{2}$, where each variable indeed carries information about $Y$ on its own. We suggest a refinement in interpreting the relationship between $X_{1}$ (or $X_{2}$) and $Y$: they are, in fact, independent, and thus it may not be appropriate to discuss a direct causal effect in this context. At the same time, representing the system by an empty graph and denying causal status to both variables seems incorrect as well, as that would collide with the case of completely unpredictable $Y$. Thus, as only the multivariate information $I(\{X_{1},X_{2}\},Y)$ is non-zero (in particular 1 bit), we suggest to state that there is no causal link, but there is a causal hyperlink from the set $\{X_{1},X_{2}\}$ to $Y$. Consequently, this perspective necessitates a shift from representing the system using directed graphs to employing directed causal hypergraphs. Note that directed hypergraph is a natural generalization of a directed graph, and corresponds to a pair $(V,E)$, where $V$ is a set of elements called nodes and $E$ is a set of pairs of subsets of $V$. Each of these pairs $(D,C)\in E$ is called a hyperedge; the vertex subset $D$ is known as its tail, and $C$ as its head. Less formally one can speak of hypernetworks with hyperlinks between (sets of) nodes, in a causal context from sources to targets. However, note that interpreting the graph-theoretical properties of ”networks” inferred from statistical dependencies is in many situations problematic and generally a challenging enterprise; see [6] as an example.
@@
While the XOR function may seem artificial, higher-order interactions are commonly discussed in the context of modelling complex systems in physics [7], neuroscience [8, 9], ecology [10], social sciences [11] and many more. Despite this challenge is increasingly recognized in the causal inference community [12, 13, 14], formal refinement of multivariate causality is not yet fully resolved.
@@ II Transfer entropy & Problem statement
In the following, we assume that for each target variable $Y$, a set of candidate source variables $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ is known a priori. In practice, the assumption is justified either by Granger causality-like reasoning: the candidate source variables are those temporally preceding the target variable, or by (additional) theoretical arguments. The transfer entropy (the positivity of which defines the presence of causal effect from $X_{i}$ to $Y$) is defined as the conditional mutual information between $X_{i}$ and $Y$ conditioned on all other potential sources:
@@
As already mentioned, we aim to generalize the transfer entropy concept to more faithfully represent causal structure in the presence of higher-order causal interactions. A natural generalization is to define existence of a causal effect of a set $\mathbf{X}_{I}=\{X_{i_{1}},\ldots,X_{i_{k}}\}$ on target $Y$ as:
@@
What would such generalization of TE conclude? First, it infers a causal link from the set $\{X_{1},X_{2}\}$: $I(\{X_{1},X_{2}\},Y)=1$ bit. However, it also infers causal link from $X_{1}$ (or $X_{2}$), as:
@@
As a side note, in practice, these individual links might remain undetected, as commonly used causal network inference algorithms progress iteratively from testing univariate predictors and may stop before discovering such conditional dependence – see [13] for discussion, and [14] for an example of an updated (yet, therefore, slower) algorithm more robust in this regard. Also note that the XOR example is straightforwardly generalizable to the k-variable higher order interaction by considering e.g. $Y$ as a sum of variables $X_{1},X_{2},\ldots X_{k-1}\stackrel{{\scriptstyle iid}}{{\sim}}Be(0.5)$ modulo 2, i.e. the parity function. Again, even the knowledge of all but one of the source variables provides no information about $Y$ without knowing the last one; it is the interaction of all $k-1$ sources that provides the target predictability.
Setting aside the implementation details, even such multivariate TE definition does not distinguish the distinctive synergistic causal effect in the XOR example, compared to, e.g. the case of summing two independent normally distributed variables. Two issues arise: first, that TE concludes that $X_{1}$ has a ’causal effect’ on $Y$ although $X_{1}$ does not hold any information about $Y,$ might be considered superfluous, as the information ’belongs’ to the pair (which multivariate TE correctly marks). Second, for the sum function, as the pair does not carry information on top of each variable, one might not want to mark the pair as causal per se, but only each variable.
@@ III Causal effect
The system is constructed (see Fig. 1) such that $X_{1}$ does not directly affect $Y$ by itself, but it affects $Y$ indirectly through mediating variables $X_{p+1},\ldots,X_{n}.$ Also $X_{1}$ affects $Y$ through pairwise interactions: each pair $\{X_{1},X_{i}\}$ for $i\in\{2,\ldots,p\}$ affects $Y$. Note that both the mutual information $I(X_{1},Y)$ and the fully conditional mutual information (transfer entropy) $I\left(\left(X_{1},Y\right)|\mathbf{X}\smallsetminus\{X_{1}\}\right)$ are greater than zero, suggesting thus a causal effect of $X_{1}$ on $Y$. In fact, all of the conditional mutual informations $I(X_{1},Y|\mathbf{S})$ are positive, except $I\left(X_{1},Y|X_{p+1},\ldots,X_{n}\right)$.
Figure 1: Example system (5) illustrating that careful conditioning is necessary to uncover that $X_{1}$ has only mediated and joint multivariate causal effects on $Y$. Linear terms are represented by oriented edges from a source element to the target element (the element on the left-hand side of the equation), while XOR terms are represented by oriented (hyper)edges from a pair of source elements to the target element.
@@
Definition (Unaccountable causal effect & Optimally conditioned causal entropy).
Let $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ be a set of candidate source variables and $Y$ a target variable. There is a causal effect from set $\mathbf{X}_{I}=\{X_{i_{1}},\ldots,X_{i_{k}}\}$ to $Y$ if and only if:
@@
In words, the causal effect is the minimum information a variable carries about a target, across conditioning it by all possible subsets of the ’rest of the universe’ - instead of considering only exactly the ’rest of the universe’ in the condition. Informally, a variable is only considered truly causal if it can not be conditioned out, whichever subset of other variables we choose.
@@
Notably, the definition of the causal effect of a set of variables typically enforces the causal influence of an arbitrary set containing a variable with a causal effect. Consider, for example, a simple linear system where the variable $Y$ is causally affected just by variable $X_{1}$ plus some noise: $Y=X_{1}+\mathcal{E}_{Y}$, and in the system, there are also other variables $\{X_{2},\ldots,X_{n}\}$ independent on $Y$ and $X_{1}$. Then, all sets containing $X_{1}$ are also causally affecting $Y$. Note that similar behaviour (supersets of causes defined as causes irrespective of further value in prediction) also holds for (a multivariate variant of) transfer entropy. In other words, although the definition of causal effect avoids attributing higher-order causal effects to individual variables, it still does not by itself distinguish multivariate causes that are trivial in the sense of inherited from its subsets from nontrivial higher-order multivariate causal effects.
We suggest one may add this distinction of a unique higher-order causal effect by interpreting differently the observation of multivariate causes in the presence/absence of the causal effects of the subsets. Further, to distinguish purely additive from synergistic multivariate causality, one may quantify the higher order causal information not induced trivially by the subsets; suitable approaches may include maximum entropy definition of higher order interactions [8, 15], or partial information decomposition (PID) [16]. The relationship between OCTE and PID is inherently close but non-trivial. For three variables, PID can be constructed using OCTE: the unique information of an individual variable can be defined as its OCTE, while the redundancy and synergy elements are derived from the multivariate information. This construction satisfies the axioms of PID. However, a more natural approach, defining unique information and synergy directly as OCTE (with synergy corresponding to the OCTE of the set of source variables)—fails to meet the PID axioms. A more sophisticated approach to PID, incorporating the strategy of powerset minimization as utilized in the analysis of how environmental factors influence transfer entropy between two variables, was recently proposed in  [17].
@@
Returning to the XOR example, note that the resulting causal graph of the system also depends on the distribution of the variables $X_{1}$ and $X_{2}.$ Consider three different distributions, see Table 1. In all cases $p_{A},p_{B},p_{C}$, $X_{1}$ and $X_{2}$ are independent variables with Bernoulli distribution.
$a)$ $b)$ $c)$
$p(X_{1}=0,X_{2}=0,Y=0)$ 0.25 0.10 0.06
$p(X_{1}=0,X_{2}=1,Y=1)$ 0.25 0.40 0.24
$p(X_{1}=1,X_{2}=0,Y=1)$ 0.25 0.10 0.14
$p(X_{1}=1,X_{2}=1,Y=0)$ 0.25 0.40 0.56
$I(X_{1},Y)$ 0.00 0.28 0.24
$I(X_{2},Y)$ 0.00 0.00 0.08
$I(X_{1},Y|X_{2})$ 1.00 1.00 0.88
$I(X_{2},Y|X_{1})$ 1.00 0.72 0.72
$I(\{X_{1},X_{2}\},Y)$ 1.00 1.00 0.96
Table 1: Example infromation-theoretical functionals for a triplet of variables coupled by the function $XOR(X_{1},X_{2}),$ where $X_{1}$ and $X_{2}$ are independent random variables with distributions: $i)$ $p_{A}:\ X_{1},X_{2}\sim Be(0.5),$ $ii)$ $p_{B}:\ X_{1}\sim Be(0.5),\ X_{2}\sim Be(0.8)$ $iii)$ $p_{C}:\ X_{1}\sim Be(0.7),\ X_{2}\sim Be(0.8).$ Note that the presence of non-zero (conditional) information values depends not only on the coupling function, but substantially on the input variable distributions, even under independence of inputs.
@@
We have already discussed the case A with uniformly distributed probabilities. In case B, $X_{1}$ is uniformly distributed, but $X_{2}\sim Be(0.8).$ While $I(X_{2},Y)$ remains zero, $I(X_{1},Y)$ is positive, as is $I(X_{1},Y|X_{2}).$ Variable $X_{1}$ thus has a causal effect on $Y$ only due to the change in the distribution $X_{2}.$ The further the variable $X_{2}$ deviates from the uniform distribution, the more information $X_{1}$ has about $Y$ and the unique contribution of the pair $\{X_{1},X_{2}\}$ decreases. In case C, where both variables deviate from the uniform distribution, both $X_{1}$ and $X_{2}$ have a causal effect on $Y$, and they also have a unique contribution as a pair $\{X_{1},X_{2}\}.$
It is worth mentioning that a similar challenge in defining causal effects, as seen in examples such as the XOR function, also arises in Judea Pearl’s approach to causality [18]. Pearl’s total effect of a variable $X$ on an outcome $Y$ is defined as a difference in the probability distribution of $Y$ when $X$ is manipulated versus when $X$ is observed naturally. Specifically, the total effect is determined by comparing the probabilities $P(Y_{x}=y)$ and $P(Y=y)$. Here, $P(Y_{x}=y)$ denotes the probability that $Y$ takes the value $y$ after an intervention on $X$ (i.e., when $X$ is set to a specific value), while $P(Y=y)$ represents the probability that $Y$ takes the value $y$ under unperturbed, observational conditions (without any intervention on $X$). If $P(Y_{x}=y)\neq P(Y=y)$, this inequality indicates that the intervention on $X$ has caused a shift in the distribution of $Y$, suggesting the existence of a total causal effect.
Further, the direct causal effect of $X$ on $Y$, holding other variables $\mathbf{Z}$ constant, is defined as the difference $P(Y_{x\mathbf{z}}=y)-P(Y_{x^{*}\mathbf{z}}=y)$, where $Y_{x\mathbf{z}}$ represents the outcome after both $X$ and $\mathbf{Z}$ are set to specific values ($X=x$ and $\mathbf{Z}=\mathbf{z}$), and $Y_{x^{*}\mathbf{z}}$ represents the outcome after $X$ is set to $x^{*}$ (different from $x$) while $\mathbf{Z}$ are set equal to the previous case ($\mathbf{Z}=z$[19]. This definition aims to isolate the direct effect of $X$ on $Y$, removing the influence of other variables.
When applying the concept of total causal effect to the example $Y=XOR(X,Z)$, we observe that $P(Y_{x}=y)=\text{Be}(0.5)=P(Y=y)$ for any value of $x$, interpreted as the absence of total causal effect of $X$ on $Y$. However, $Y_{x=0,z=0}=0$ and $Y_{x=1,z=0}=1$, therefore indicating a direct effect of $X$ on $Y$.
The absence of total causal effect in the presence of a direct causal effect appears contradictory and difficult to interpret. However, we propose one can apply the same strategy as in the case of transfer entropy, in particular by redefining a variable as having a direct causal effect if and only if its perturbations affect the distribution irrespective of which subset of the other system variables is kept fixed.